### v1

@@ 3.1 Characteristic modes of crack growth in irradiated structures
unirradiated sample
sample irradiated to 0.152 dpa
sample irradiated to 0.266 dpa
@@
In the unirradiated sample (top row of Figs 5 and 7), the crack advances with minimal resistance, exhibiting a relatively smooth growth path accompanied by a developed plastic zone at the crack tip. The emission of dislocations near the crack tip is observed to hinder crack propagation by redistributing stress and promoting localized plastic deformation. This dislocation activity contributes to crack tip blunting (Fig. 7: at 75 to 100 ps), as evidenced by the gradual widening of the crack tip and the formation of a diffuse plastic zone (Fig. 5: at 70 to 100 ps).
In the irradiated samples, the crack growth dynamics is distinctly different. The sample with 0.152 dpa (Fig. 7: middle row) shows a crack path that is restricted and/or redirected under the influence of radiation-induced defects. The crack path is altered by the presence of radiation-induced defects, which act as local strain concentrators. These defects intensify dislocation emission and change the crack propagation behavior. Unlike the unirradiated sample in Fig. 5, where dislocations are primarily emitted from the crack tip, the irradiated sample exhibits additional dislocation activity originating from radiation-induced defects (Fig. 5, sample irradiated to 0.152 dpa). These defects not only concentrate strain and contribute to higher dislocation density but also limit the mobility of the emitted dislocations, thereby restricting plastic deformation and modifying the crack’s trajectory. As a result, an amorphous zone with a varying width of 2–5 nm is formed at the crack tip (in Fig. 7, marked with a pink frame). The zone evolves dynamically, propagating together with the crack tip.
Furthermore, under mechanical loads, as the crack advances, radiation-induced voids expand and coalesce, forming larger void structures. The blue frames in Fig. 7 highlight instances where previously isolated voids join together. This coalescence is driven by localized plastic deformation, which facilitates void merging and provides an easier path for crack propagation. Notably, this mechanism is analogous to void coalescence observed in non-irradiated ductile materials [42]. Near the crack tip, coalescing voids create additional weak zones, directing the crack toward areas of lower resistance. This process enhances crack branching and influences its overall morphology. Eventually, as the crack encounters clusters of coalesced voids, these merged structures integrate into the main crack, leading to more irregular and complex crack propagation patterns (Fig. 7; at 100 ps).
@@
Fig. 7 reveals the dynamic interaction between crack progression and void evolution, demonstrating how void growth, void coalescence, and subsequent merging with the crack enhance crack propagation. Moreover, Fig. 7 discloses that in irradiated materials, secondary cracks can form if the stress level in the sample is sufficiently high and the primary crack is blocked by tangled dislocations or amorphous zones. When radiation voids develop and coalesce in other regions of the material, these zones become weakened and may serve as initiation sites for secondary cracks. This process ultimately leads to a complex fracture morphology that is significantly influenced by void accumulation and absorption.
sample irradiated to 0.07 dpa
          t=5ps                            t=10ps                            t=30ps                            t=40ps
Refer to captionRefer to captionRefer to captionRefer to caption
sample irradiated to 0.494 dpa
          t=5ps                            t=10ps                            t=25ps                            t=40ps
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 8: Interaction between the crack tip and radiation-induced defects. The colors representing different defect types are explained in the legend of Fig. 5.
Fig. 8 shows the interaction between the crack front and radiation defects at the early stage. In regions with high void concentrations (Fig. 8; sample irradiated to 0.07 dpa; top row), cracks can propagate more easily as the local lattice structure is weakened. The presence of radiation voids lowers effectively the cohesive energy, making the material more susceptible to crack initiation under mechanical loading [43]. As the crack propagates through regions with high concentrations of vacancy clusters, its growth is accelerated due to the reduced local resistance to fracture (Fig. 8; sample irradiated to 0.07 dpa; top row). The crack tip grows through the space previously occupied by the void.
Another critical mechanism that influences crack growth and changes the material’s fracture behavior is the interaction between a propagating crack and radiation-induced dislocation loops (Fig. 8; sample irradiated to 0.494 dpa). The radiation-induced dislocation loops act as barriers to dislocation motion, reducing dislocation mobility near the crack tip. When the crack encounters a dislocation loop, the loop obstructs further dislocation propagation and creates local zones of high resistance (Fig. 8; sample irradiated to 0.494 dpa; at 40 ps).
These mechanisms analysis highlight how voids and loops exert contrasting effects on crack behavior, with voids accelerating propagation and dislocation loops impeding it. For instance, radiation-induced dislocation loops can cluster near the crack front, forming a knot of entangled dislocations that further complicates the crack propagation process. In this case, the crack’s progress is slowed or halted as it encounters this dense, immobile cluster of dislocations. This mechanism is shown in Fig. 9 for different values of dpa at selected time instants. This structure acts as a strong inclusion that prevents further crack advancement, effectively blocking the crack path. The accumulated loops at the crack tip serve as a physical barrier that limits crack propagation by restricting dislocation mobility. In higher-dpa samples, secondary cracks can form as they initiate on the radiation defects, where the critical energy for atomic bond breaking is reached due to stress concentrations (see Fig. 7). This behavior illustrates a transition toward crack branching, where the main crack’s path is impeded, leading to the initiation and coalescence of secondary cracks that eventually merge with the main crack. This phenomenon, driven by high defect density, can contribute to the complex fracture patterns seen in highly irradiated materials [44, 45, 46].
dpa=0.07; t=95ps                          dpa=0.076; t=115ps                     dpa=0.494; t=100ps Refer to caption Refer to caption Refer to caption
@@ 3.2 Fundamental mechanisms governing defect–crack interactions
Figure 10: Fundamental mechanisms of a single defect–crack interaction governing crack propagation. Frames outlined in pink correspond to the same moment in time (75ps).
The interaction between radiation-induced defects and propagating cracks plays a pivotal role in defining the fracture behavior of irradiated materials. While previous sections explored the broad impact of irradiation on crack propagation dynamics and the microstructural changes induced by defects, this subsection focuses on analyzing specific, fundamental interactions between individual defects and a crack front.
Fig. 10 captures representative cases of defect–crack interactions, showcasing how different defect types-small voids, large voids, dislocation loops, and rigid inclusions-influence crack morphology and behavior. Each row of Fig. 10 corresponds to a distinct scenario, selected for its ability to illustrate the diversity of interactions and their implications for crack propagation. By isolating these mechanisms, the physical principles governing crack evolution under varying microstructural conditions and irradiation levels are revealed.
The analysis highlights how certain defects, such as voids, act as facilitators for crack growth by reducing effective cohesive energy, while others, like dislocation loops and rigid inclusions, serve as barriers that redirect or even halt crack progression. These analysis provides a deeper understanding of how microstructural features influence crack dynamics and fracture patterns.
@@
In the third row of Fig. 10, the crack interacts with a dislocation loop spanning 5 nm. This interaction is distinct because the dislocation loop acts as a barrier, impeding crack growth by obstructing dislocation motion. As the crack encounters the dislocation loop, it experiences increased resistance, which can alter its trajectory or even temporarily halt its progression. The presence of the dislocation loop leads to localized stress concentrations around the crack front, which can enhance dislocation emission and create complex dislocation structures near the crack tip.
In the fourth row of Fig. 10, the crack faces two rigid inclusions, each with dimensions of 5 nm. The rigid inclusions are introduced through the application of boundary conditions that constrain any deformation within the inclusion region during the simulation, thereby ensuring the inclusion remains a fixed, immovable obstacle within the material. These inclusions represent strong obstacles that divert the crack path, leading to a more tortuous and irregular morphology. Unlike voids, which weaken the material locally, rigid inclusions increase local resistance and force the crack to change direction, often leading to stress concentration at points where the crack path is redirected. This can result in additional dislocation activity around the inclusions as the crack adapts to the path of least resistance. This results in a more tortuous and branched crack morphology.
Fig. 10 presents the interplay between crack dynamics and radiation-induced defects, revealing how defect size, type, and spatial distribution shape crack behavior. This analysis emphesizes the dual role of defects in either facilitating or obstructing crack propagation, depending on their nature. The examples examined highlight the transition from crack paths governed by localized weakening, as seen with voids, to trajectories shaped by resistance and redirection due to dislocation loops or rigid inclusions.
@@ 3.3 Summary of defect-driven fracture mechanisms in irradiated materials
  • Mechanism of void–crack interactions: Voids act as local weak points in the material, reducing the effective energy and creating pathways of reduced resistance for crack growth. In regions with high void concentrations, the crack propagates more easily, as the voids lower the local stress required for atomic bond breaking. This mechanism is particularly pronounced when voids coalesce ahead of the crack tip, forming larger voids that further weaken the material. The coalescence of voids not only accelerates crack propagation but also leads to irregular crack morphologies, as the crack tends to follow the path of least resistance through these weakened zones. This process often results in crack branching, where the primary crack splits into multiple secondary cracks, increasing the overall fracture surface area and contributing to a more complex fracture pattern.
@@
  • Mechanism of dislocation loop–crack interactions: In contrast to voids, dislocation loops act as barriers to crack propagation. These loops, formed from clusters of interstitial atoms that form stable structures, obstruct the motion of dislocations near the crack tip, creating localized stress concentrations. When the crack encounters a dislocation loop, its progression is temporarily halted. This interaction can lead to the emission of additional dislocations from the crack tip, promoting localized plastic deformation and energy dissipation. However, in irradiated materials, the mobility of these dislocations is significantly restricted by the presence of other radiation-induced defects, such as voids and additional dislocation loops. As a result, the energy dissipation through plasticity is reduced, and the material becomes more prone to brittle fracture. Notably, dislocation loops, due to their immobile nature, can also act similarly to rigid inclusions by deflecting crack paths, increasing local stress concentrations, and altering crack propagation trajectories.
@@
  • Radiation-induced defects as nucleation sites for dislocations: In addition to acting as obstacles, irradiation-induced defects serve as preferential nucleation sites for dislocation formation. Defect clusters, such as voids and dislocation loops, introduce local stress concentrations that facilitate dislocation emission. As irradiation levels increase, dislocation nucleation leads to reduced plasticity and a more brittle fracture response. This dual role of defects, both as dislocation sources and barriers, critically influences the balance between ductile and brittle failure mechanisms.
@@
  • Mechanism of amorphous zone development at the crack tip: The amorphous zones which consist of disordered atomic structures form as a result of the accumulation of radiation-induced defects, such as dislocation loops. These zones, act as barriers to crack propagation, absorbing energy and temporarily halting the crack’s advance. However, these zones also contribute to localized embrittlement, as the disordered structure reduces the material’s ability to deform plastically. As a result, the secondary cracks can be formed, further increasing the complexity of the fracture process.
@@
  • Crack branching and secondary crack formation: High defect densities, particularly clusters of dislocation loops, lead to crack branching and the initiation of secondary cracks. When the primary crack encounters dense defect fields, it can branch or form secondary cracks, which eventually coalesce with the main crack. This results in irregular fracture patterns and an increased fracture surface area. The formation of secondary cracks is driven by the stress concentrations around radiation-induced defects, which provide initiation sites for new cracks.
@@ 4.1 Framework for energy-based fracture analysis
Recent advances by Xu and Demkowicz [20] have applied this framework to atomistic simulations, directly calculating the critical energy release rate $G_{\rm c}$ for systems with crack-tip plasticity. It has been shown that $w_{\rm p}$ depends on the sample size, which is strongly limited in MD due to computational cost. In this approach, $G_{\rm c}$ is defined as the increment of the total dissipated energy, $\mathrm{d}E_{\rm dis}$, per increment of the projected crack surface area, $\mathrm{d}A_{\rm crack}$, thus
@@
Here, referring to the MD simulations, the dissipated energy increment $\mathrm{d}E_{\rm dis}\!=\!\mathrm{d}W\!-\mathrm{d}U$ is determined as the difference between the increments of the internal energy $U$ (comprising potential and kinetic energies) and external work $W$ (derived from the stress–strain response). Note that, under isothermal conditions, the dissipated energy corresponds to the change in Helmholtz free energy, $\mathrm{d}E_{\rm dis}=\mathrm{d}W\!-\mathrm{d}F$ were $\mathrm{d}F=\mathrm{d}U-T\mathrm{d}S$, which includes contributions from both internal energy $U$ and entropy $S$. Above, the entropy contribution has been disregarded.
Moreover, for processes maintained at constant temperature, which define the scope of this analysis, the total dissipated energy reflects contributions from both bond rupture and plastic deformation. In the MD framework applied here, the thermostat regulates temperature by dynamically adjusting atomic velocities to dissipate excess kinetic energy generated during plastic work. This ensures the system remains at constant temperature ($T$=constant), isolating mechanical contributions to $\mathrm{d}E_{\rm dis}=\mathrm{d}W\!-\mathrm{d}U$.
4.2 Effects of irradiation on fracture energy

### v2

@@ 3.1 Characteristic modes of crack growth in irradiated structures
In the unirradiated sample (top row of Figs 4 and 5), the crack advances with minimal resistance, exhibiting a relatively smooth growth path accompanied by a developed plastic zone at the crack tip. The emission of dislocations near the crack tip is observed to hinder crack propagation by promoting localized plastic deformation. This dislocation activity contributes to crack tip blunting (Fig. 5: at 75 to 100 ps), as evidenced by the gradual widening of the crack tip and the formation of a diffuse plastic zone (Fig. 4: at 70 to 100 ps).
In the irradiated samples, the crack growth dynamics is distinctly different. The sample with 0.152 dpa (Fig. 5: middle row) shows a crack path that is restricted and/or redirected under the influence of radiation-induced defects. Unlike the unirradiated sample in Fig. 4 (top row, unirradiated sample), where dislocations are primarily emitted from the crack tip, the irradiated sample exhibits additional dislocation activity originating from radiation-induced defects (Fig. 4, sample irradiated to 0.152 dpa). The defects not only concentrate strain and contribute to higher dislocation density but also limit the mobility of the emitted dislocations, thereby restricting plastic deformation and modifying the crack’s trajectory. As a result, an amorphous zone with a varying width of 2–5 nm is formed at the crack tip (in Fig. 5, marked with a pink frame). The zone evolves dynamically, propagating together with the crack tip. Furthermore, under mechanical loads, as the crack advances, radiation-induced voids expand and coalesce, forming larger void structures. The blue frames in Fig. 5 highlight instances where previously isolated voids join together. This coalescence is driven by localized plastic deformation, which facilitates void merging and provides an easier path for crack propagation. Notably, this mechanism is analogous to void coalescence observed in non-irradiated ductile materials [48]. Near the crack tip, coalescing voids create additional weak zones, directing the crack toward areas of lower resistance. This process enhances crack branching and influences its overall morphology. Eventually, as the crack encounters clusters of coalesced voids, these merged structures integrate into the main crack, leading to more irregular and complex crack propagation patterns (Fig. 5; at 100 ps).
@@
Fig. 5 reveals the dynamic interaction between crack progression and void evolution, demonstrating how void growth, void coalescence, and subsequent merging with the crack enhance crack propagation. Moreover, Fig. 5 discloses that in irradiated materials, secondary cracks can form if the stress level in the sample is sufficiently high and the primary crack is blocked by tangled dislocations or amorphous zones [22]. When radiation voids develop and coalesce in other regions of the material, these zones become weakened and may serve as initiation sites for secondary cracks. This process ultimately leads to a complex fracture morphology that is significantly influenced by void accumulation and absorption.
@@
Another critical mechanism that influences crack growth and alters fracture behavior is the interaction between the propagating crack and radiation-induced dislocation loops. This mechanism is illustrated in Fig. 6 for different dpa levels. The dislocation loops create entangled structures that act as barriers to crack advance by reducing dislocation mobility near the crack tip. When the crack front encounters these obstacles, further propagation is hindered, and localized zones of increased resistance to fracture develop.
The above analysis highlights the contrasting roles of voids and dislocation loops in crack propagation. Voids facilitate crack advance by reducing local cohesion, while dislocation loops act as barriers that slow down or block the crack front. At higher dpa levels, these effects can coexist—crack branching may occur when the main crack is obstructed by loop accumulations, while secondary cracks initiate in weakened regions around radiation voids (see Figs. 5 and  6). Together, these competing mechanisms contribute to the complex fracture morphology observed in irradiated materials.
@@ 3.2 Fundamental mechanisms governing defect–crack interactions
The previous subsection examined the overall effect of irradiation on crack dynamics and microstructure evolution. This section focuses on selected examples of how isolated, typical radiation-induced defects interact with the crack front.
Fig. 7 presents representative atomistic simulations showing the effect of specific defect types, small and large voids, dislocation loops, and rigid inclusions on crack propagation. Rigid inclusions are used here as a simplified model to mimic the strong barrier effect of highly entangled dislocation structures. Each case illustrates a distinct interaction mechanism and its influence on the crack path and morphology. This approach allows to isolate and better understand the fundamental defect–crack interaction mechanisms that contribute to the macroscopic fracture behavior of irradiated alloys.
@@
Figure 7: Fundamental mechanisms of a single defect–crack interaction governing crack propagation. The individual rows correspond to small voids (top), large voids, dislocation loop, and rigid inclusion (bottom). Frames outlined in pink correspond to the same moment in time (75ps).
@@
In the third row of Fig. 7, the crack interacts with a dislocation loop spanning 5 nm. This interaction is distinct because the dislocation loop acts as a barrier, impeding crack growth by obstructing dislocation motion. As the crack encounters the dislocation loop, it experiences increased resistance, which can alter its trajectory or even temporarily halt its progression.
In the fourth row of Fig. 7, the crack faces two rigid inclusions, each with dimensions of 5 nm. The rigid inclusions are introduced through the application of boundary conditions that constrain any deformation within the inclusion region during the simulation, thereby ensuring the inclusion remains a fixed, immovable obstacle within the material. These inclusions represent strong obstacles that divert the crack path, leading to a more tortuous and irregular morphology. Unlike voids, which weaken the material locally, rigid inclusions increase local resistance and force the crack to change direction. This can result in additional dislocation activity around the inclusions as the crack adapts to the path of least resistance. This results in a more tortuous and branched crack morphology.
Fig. 7 shows how the crack response depends on defect type and size. This analysis confirms that voids act as local weakening sites that promote crack growth, while dislocation loops and rigid inclusions impede propagation by increasing resistance near the crack tip. The transition from facilitation to obstruction highlights the complex role of microstructural features in shaping crack morphology.
These insights complement the broader findings from irradiated samples discussed in the previous subsection. While Section 3.1 explored collective crack behavior across varying irradiation levels, the current analysis in Section 3.2 isolates individual mechanisms to reveal their distinct contributions to fracture evolution in irradiated alloys.
@@ 3.3 Summary of defect-driven fracture mechanisms in irradiated materials
  • Mechanism of void–crack interactions:  Voids act as local weak points in the material, reducing the effective energy and creating pathways of reduced resistance for crack growth. In regions with high void concentrations, the crack propagates more easily, as the voids lower the local stress required for atomic bond breaking. This mechanism is particularly pronounced when voids coalesce ahead of the crack tip, forming larger voids that further weaken the material. The coalescence of voids not only accelerates crack propagation but also leads to irregular crack morphologies, as the crack tends to follow the path of least resistance through these weakened zones. This process often results in crack branching, where the primary crack splits into multiple secondary cracks, increasing the overall fracture surface area and contributing to a more complex fracture pattern. Crack–void interaction mechanisms with artificially introduced defects were previously studied using molecular dynamics in single-crystal silicon [26], where the effect of void spacing was analyzed. Similar simulations in aluminum [50] showed that vacancy clusters can alter crack tip behavior by modifying stress concentrations and promoting crack tip blunting.
@@
  • Mechanism of dislocation loop–crack interactions:  In contrast to voids, dislocation loops act as barriers to crack propagation. These loops, formed from clusters of interstitial atoms that form stable structures, obstruct the motion of dislocations near the crack tip, creating localized stress concentrations. When the crack encounters a dislocation loop, its progression is temporarily halted. This interaction can lead to the emission of additional dislocations from the crack tip, promoting localized plastic deformation and energy dissipation. However, in irradiated materials, the mobility of these dislocations is significantly restricted by the presence of other radiation-induced defects, such as voids and additional dislocation loops. As a result, the energy dissipation through plasticity is reduced, and the material becomes more prone to brittle fracture. Notably, dislocation loops, due to their immobile nature, can also act similarly to rigid inclusions by deflecting crack paths, increasing local stress concentrations, and altering crack propagation trajectories. Dislocation–crack interactions have also been investigated in earlier atomistic studies on ideal crystalline systems [52, 53]. MD simulations in nickel showed that dislocations approaching the crack tip can trigger dislocation emission, cross-slip, and the development of a localized plastic zone [49]. However, these studies did not consider radiation-induced dislocation loops or their evolution, and therefore do not capture the complex mechanisms present in irradiated materials.
@@
  • Radiation-induced defects as nucleation sites for dislocations:  In addition to acting as obstacles, irradiation-induced defects serve as preferential nucleation sites for dislocation formation. Defect clusters, such as voids and dislocation loops, introduce local stress concentrations that facilitate dislocation emission. As irradiation levels increase, dislocation nucleation leads to reduced plasticity and a more brittle fracture response. This dual role of defects, both as dislocation sources and barriers, critically influences the balance between ductile and brittle failure mechanisms. This behavior was also observed in our study through atomistic simulations of shear deformation in irradiated Fe–Ni–Cr alloys, where dislocation nucleation was observed at radiation-induced defect clusters [14].
@@
  • Mechanism of amorphous zone development at the crack tip:  The amorphous zones which consist of disordered atomic structures form as a result of the accumulation of radiation-induced defects, such as dislocation loops. These zones, act as barriers to crack propagation, absorbing energy and temporarily halting the crack’s advance. However, these zones also contribute to localized embrittlement, as the disordered structure reduces the material’s ability to deform plastically. As a result, the secondary cracks can be formed, further increasing the complexity of the fracture process. Similar amorphization-induced crack shielding has also been reported in Si single crystals under high contact stresses during grinding, where amorphous zones were shown to suppress cracking by altering the local stress state [22].
@@
  • Crack branching and secondary crack formation:  High defect densities, particularly clusters of dislocation loops, lead to crack branching and the initiation of secondary cracks. When the primary crack encounters dense defect fields, it can branch or form secondary cracks, which eventually coalesce with the main crack. This results in irregular fracture patterns and an increased fracture surface area. The formation of secondary cracks is driven by the stress concentrations around radiation-induced defects, which provide initiation sites for new cracks. Crack branching and secondary cracks are often observed in MD simulations for brittle materials, e.g. [54, 55], but also for ductile materials, e.g., for titanium under cyclic loading [56].
@@ 4 Analysis of the critical strain energy release rate
This section establishes a theoretical and computational framework for determiening the critical strain energy release rate ($G_{\rm c}$) in irradiated materials, bridging classical fracture mechanics with atomistic simulations. It examines the interplay between bond rupture (surface energy, $\gamma_{\rm s}$) and dislocation-mediated plasticity (plastic work, $w_{\rm p}$) under irradiation, using molecular dynamics to isolate irradiation- and temperature-dependent contributions to $G_{\rm c}$.
@@ 4.1 Framework for energy-based fracture analysis
Recent advances by Xu and Demkowicz [21], see also Barik et al. [9], have applied this framework to atomistic simulations, directly calculating the critical energy release rate $G_{\rm c}$ for systems with crack-tip plasticity. It has been shown that $w_{\rm p}$ depends on the sample size, which is strongly limited in MD due to computational cost. In this approach, $G_{\rm c}$ is defined as the increment of the total dissipated energy, $\mathrm{d}E_{\rm dis}$, per increment of the projected crack surface area, $\mathrm{d}A_{\rm crack}$, thus
@@
Here, referring to the MD simulations, the dissipated energy increment $\mathrm{d}E_{\rm dis}\approx\mathrm{d}W-\mathrm{d}U$ is determined as the difference between the increments of the external work $W$ (derived from the stress–strain response) and internal energy $U$ (comprising potential and kinetic energies). Note that, under isothermal conditions (maintained by the thermostat), the dissipated energy corresponds to the change in Helmholtz free energy, $\mathrm{d}E_{\rm dis}=\mathrm{d}W\!-\mathrm{d}F$, where $\mathrm{d}F=\mathrm{d}U-T\mathrm{d}S$, which includes contributions from both internal energy $U$ and entropy $S$. Since the entropy and its changes cannot be computed for the problem at hand, the entropy contribution is disregarded above so that $\mathrm{d}U$ is assumed to approximate $\mathrm{d}F$, thus $\mathrm{d}F\approx\mathrm{d}U$, as tacitly assumed in [21, 9]. As a result, Eq. (2) is obtained, and thus the heat generated during the inelastic deformation process (and taken from the system by the thermostat to maintain a constant temperature) is adopted as a measure of the dissipated energy. Note that it is commonly accepted that the majority of plastic work is converted to heat and thus dissipated (the rest being stored in the material), although in the context of MD simulations this depends on the material, the stage of the deformation process, and whether the integral or instantaneous (differential) energy conversion rate is considered [60, 61]. Nevertheless, the generated heat constitutes the main contribution in the energy balance, which justifies the above approximation.
4.2 Effects of radiation on fracture energy